OpenStudy (anonymous):

Can some help me. I'm totally stumped with this problem: y varies inversely with x, and x = 120 when y =90. Find the inverse variation equation and use it to find the value of x when y = 30. 210 40 360 22.5 y varies inversely with x. Find the inverse variation equation if x = 1.4 when y = 70. I got the first part of the answer xy = 98 y/x = 98 The problem is that I'm not sure how the xy goes. Thank you

OpenStudy (cwrw238):

inverse variation is y = k/x where k is a constant x = 120 when y = 90 90 = k/120 k = 10800 so relation between y and x is y = 10800/x when y = 30 30 = 10800 / x x = 10800 / 30 = 360